Apr 9, 2015 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Work with Hovanes to activate VME LED pulser setup for FCAL (Elton). Note: Currently the system still works using the associated NIM electronics. Hovanes has suggested that we delay the new programing till after this run so as not impact the working system. We will try to see if Orlando can devote some time to this project in late May. Otherwise, we will try to update the program ourselves.


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. Editable Wish-list of action items: BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Spring Run
  4. FCAL Update [2]
  5. BCAL Update
    1. BCAL - Layer 4 (George)
  6. Reconstruction and Simulation
  7. Calibration
  8. Any other business


Attendees: Elton, Mark, Tegan, Adesh, Will, Manuel (JLab); Zisis, Andrei, Ahmed (UofR); George, Christina (Athens)

  1. Announcements
  2. Action Items
  3. Spring Run
    1. FCAL
      • No updates on hardware. Only one base is currently off, which occurred when one IOC was rebooted.
      • Manuel has checked that the addition of the FCAL parts works as expected in the test JInventory DB. Paul Smith (IU) had some reservations about the mechanics of operation, but is willing to use it for tracking problems. During a phone conference on Tuesday (Manuel, Sergey, David and Elton) it was decided that Manuel could proceed with inputing all information into the working JInventory DB. He will make it available to Paul Smith when ready.
    2. BCAL
      • No updates. Biases are on 1.4 V overvoltage, chiller at 10 deg. Stable
    3. Commissioning Time Chart
      • Discussion on current schedule and priorities
      • Beam expected (hoped) for next Thu. No extension beyond May 4. That leaves about 18 days on the floor x 0.5 = 9 days at 100%.
      • Schedule will be revisited on Monday at the RC meeting
      • Zisis: Probably tgt commissioning is not top priority
      • Zisis: Develop one plan for steady beam and alternate for unsteady beam
      • Elton: FCAL should consider efficiency of pi0 collection with improved trigger, but perhaps not optimal beam
      • Adesh: He and Sasha have investigated trigger thresholds and determined that single block threshold of 160 counts substantially reduced rate (we ran with 150 last fall).
  4. FCAL Update [3]
    1. Relative estimate of pi0 rates for E=5.5 GeV relative to E=10 GeV. (Adesh)
      • Followed example by Will's analysis for BCAL
      • Used DBeamPhoton class for access to tagger information. This combines hodoscope and microscope information
      • Finds prominent pi0 peak in Fcal trig data, but relatively small peak in Fcal-Bcal trigger.
      • Conclusion is that only about 20-30% of pi0's found at 10 GeV remain after a photon energy cut of 5.5 GeV
      • Backgrounds are highly trigger dependent. Pi0 width is about 40 MeV.
  5. BCAL Update
    1. BCAL - Layer 4 (George)
      • Checked fraction of energy dumped into layer 4 with parametric shower profile and plots from GlueX-doc-1871 at 1 and 2 GeV at 90 deg
      • Somewhat less than 8.4 % of energy is deposited in layer 4 at 1 GeV and 12.7% at 2 GeV. That corresponds to 43 MeV deposited in row 7 at 1 GeV.
      • Elton: This should be above the readout threshold.
      • Zisis: Will collect simulation info and summarize for group
      • Christina: Suggest to take data with lower threshold this spring. Elton: Currently we are running at 1.4 V overbias (essentially reducing the relative DAQ threshold) and we also plan to do a threshold scan to optimize running conditions. Clearly we would like to run with as low a threshold as possible.\
      • Elton: Will could try to look at the amount of energy deposited in layer 4 with the BGGEN simulations.
  6. Reconstruction and Simulation
    1. Dark Hits (Tegan)
      • Converting MCsmear from single waveforms to single energies per cell. This is almost done.
      • Trying to adjust amplitude of dark hits to match data by looking at the pedestal widths.
      • Note after meeting: Based on Marks suggestion, decided that Will would take pedestal data scanning over overbiases in the BCAL. This could be helpful in determining contribution of dark hits to the pedestal.
    2. Etas in simulation (Will)
      • No etas are seen in the fall data. Investigating with BGGEN simulations how many we should actually expect after realistic reconstruction and selection cuts.
      • Once a signal is observed in simulation, one can try to optimize selection cuts/reconstruction to extract the eta signal.
  7. Calibration
  8. Any other business